OpenStudy (anonymous):

There are 28 students in a history class. The number of juniors is 24 less than three times the number of seniors. How many juniors and how many seniors are in the class?

OpenStudy (australopithecus):

write it out as you read it

OpenStudy (anonymous):

failed for me

OpenStudy (australopithecus):

The number of juniors is 24 less than three times the number of seniors. Can you express this?

OpenStudy (australopithecus):

times refers to multiplication, less refers to subtraction

OpenStudy (anonymous):

24x-3y=28

OpenStudy (mertsj):

What can you say about the junior and seniors?

OpenStudy (australopithecus):

Total amount of students = 28 Total amount of juniors = 24 Make up variables x and y set y equal to the number of juniors set x equal to the number of seniors for this question you will need to come up with two equations to describe this scenario then use system of equations to solve

OpenStudy (anonymous):

I hate this...

OpenStudy (mertsj):

It is so simple. Don't give up.

OpenStudy (mertsj):

Let's start over.

OpenStudy (anonymous):

Fine

OpenStudy (anonymous):

I'll try...

OpenStudy (mertsj):

How about we use j for juniors and s for seniors

OpenStudy (mertsj):

What is j+s equal to?

OpenStudy (anonymous):

j=24-3s

OpenStudy (mertsj):

Don't make it harder than it is. How many kids are in the class?

OpenStudy (anonymous):

28

OpenStudy (mertsj):

Then j+s = 28

OpenStudy (mertsj):

Now the problem says "the number of juniors is" That would be j =

OpenStudy (mertsj):

and then it says "24 less" You make something 24 less by subtracting 24. If you have $50 and I have $24 dollars less, then I have 50-24

OpenStudy (anonymous):

24-3s

OpenStudy (mertsj):

So the juniors are 24 less than three times the seniors. Well, three times the seniors is 3s and 24 less than that is 3s-24

OpenStudy (mertsj):

Did you subtract the 24?

OpenStudy (mertsj):

jojo1236 24-3s

OpenStudy (anonymous):

my answer doesn't make sense

OpenStudy (mertsj):

LOOK AT ME!!!

OpenStudy (mertsj):

Did you subtract the 24?

OpenStudy (anonymous):

OK

OpenStudy (mertsj):

Are you going to answer my question or are we done?

OpenStudy (anonymous):

I get 4

OpenStudy (anonymous):

sorry this website runs slow on my computer for some reason.idk?

OpenStudy (mertsj):

ok. j+s=24 j=3s-24 Solve the system

OpenStudy (anonymous):

what does j=?

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

does j=15 and s=13???

OpenStudy (mertsj):

yes